### v1

# Investigating Use Cases of AI-Powered Scene Description Applications for Blind and Low Vision People

Abstract.
”Scene description” applications that describe visual content in a photo are useful daily tools for blind and low vision (BLV) people. Researchers have studied their use, but they have only explored those that leverage remote sighted assistants; little is known about applications that use AI to generate their descriptions. Thus, to investigate their use cases, we conducted a two-week diary study where 16 BLV participants used an AI-powered scene description application we designed. Through their diary entries and follow-up interviews, users shared their information goals and assessments of the visual descriptions they received. We analyzed the entries and found frequent use cases, such as identifying visual features of known objects, and surprising ones, such as avoiding contact with dangerous objects. We also found users scored the descriptions relatively low on average, 2.76 out of 5 (SD=1.49) for satisfaction and 2.43 out of 4 (SD=1.16) for trust, showing that descriptions still need significant improvements to deliver satisfying and trustworthy experiences. We discuss future opportunities for AI as it becomes a more powerful accessibility tool for BLV users.
Use cases, Scene Description, AI, Blind and Low Vision People, Computer Vision, Assistive Technology, Diary Study
journalyear: 2024copyright: rightsretainedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642211isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Accessibilityccs: Human-centered computing Empirical studies in accessibilityccs: Human-centered computing Empirical studies in HCI
Figure 1. During the diary study, participants submitted entries with different goals. These examples represent four of the most common goals: (A) Getting a scene description, (B) identifying features of objects, (C) obtaining the identity of a subject in the scene, and (D) learning about the application.
18 images randomly sampled from the diary study. These examples represent four of the most common goals: The first three images are examples of (A) Getting a scene description, the next six images are examples of (B) identifying features of objects, the next four images are examples of (C) obtaining the identity of a subject in the scene, and the last five are examples of (D) learning about the application.
Figure 1. During the diary study, participants submitted entries with different goals. These examples represent four of the most common goals: (A) Getting a scene description, (B) identifying features of objects, (C) obtaining the identity of a subject in the scene, and (D) learning about the application.
1. Introduction
Blind and low vision (BLV) people face many daily challenges in mundane tasks. Seemingly simple questions can become frustrating ordeals. For example, when picking something to wear, a BLV person may wonder: “Which of these blouses is the one with pink flowers on it?” When trying to eat healthy, she might inquire: “Does this protein bar have soy?” And when looking after her son, she might need to know: “Did my toddler throw food on the living room rug?” In recent years, visual interpretation applications have become available to help BLV people answer such questions. They allow users to submit a photo or video to a remote service that returns information about its visual contents.
Researchers have studied the use of applications that describe visual information extensively. The most significant work on this topic explored the use of VizWiz and Aira, visual description applications that allow users to send a photo and question to a remote sighted assistant. Brady et al. (Brady et al., 2013) analyzed over 40,000 VizWiz submissions and categorized user questions. They found that some of the most common types of questions involved identifying objects, describing things, and reading text. Lee et al. (Lee et al., 2018) studied the use of Aira, interviewing BLV users as well as sighted assistants. Similar to Brady et al.’s findings, they found that users sought descriptions of things and asked questions to support navigation. Aira users also wanted information about social interactions. These studies shed light on use cases for visual description applications, but they only examined use of applications with remote sighted assistants. A growing number of visual description applications are powered by AI rather than humans (e.g., Seeing AI (Microsoft, 2023d) and BeMyAi (Eyes, [n. d.]))–they leverage computer vision and natural language processing rather than sighted assistants. Thus, at this time, despite the burst of interest in AI, little is known about how these AI-powered applications are being used.
Although AI and human-powered applications both address visual access needs, BLV people may use them in different ways. People’s sense of trust, perceptions of privacy, and feelings of independence are likely to differ when requesting information from AI rather than a human; this may affect when, where, and how they use an application. Prior work has found that users have both over trusted and been skeptical of AI-generated captions (Lee et al., 2021) and it is unknown how such varying levels of trust affect use of an application. Another study found that human assistance can detract from people’s sense of independence (Collins et al., 2023), which may lead to increased use of AI-powered applications compared to their human-powered counterparts.
Despite the growing prominence of AI, little work has studied the use of AI-powered scene description applications (at the time of this study, a more common deployment of AI was to describe scenes captured in images; thus, we refer to the AI-powered applications mentioned here as scene description applications). Kupferstein et al. (Kupferstein et al., 2020) briefly described a pilot diary study with four users of SeeingAI. Their preliminary findings show that participants used the application to seek private information and learn about objects they could not touch (e.g., an object behind a glass wall). Interestingly, participants were satisfied with only 43.0% of interpretations. This initial exploration raised critical questions: In what other use cases do BLV people use AI-powered scene description applications? And why were participants not satisfied with the majority of interpretations? In short, there is presently a gap in our understanding of the use of AI-powered scene description applications.
We addressed this gap by building on Kupferstein et al.’s work with a rigorous investigation of an AI-powered scene description application. We posed the following research questions:
  • What are BLV people’s use cases for AI-powered scene description applications?
  • How well are BLV people’s needs addressed in these use cases?
To answer these questions, we conducted a two-week diary study with 16 BLV participants. During this period, participants used a scene description application that we developed for this study. Working similarly to SeeingAI, the application described images using a state-of-the-art image analysis model developed by Microsoft (Microsoft, 2023a; Team, [n. d.]) and produced descriptions for photos submitted by the user. After providing a description, the application presented the participant with questions about the information they received, the user’s goal, and their use context, which all constituted a diary entry. At the end of the diary period, we conducted interviews with participants where we reviewed their entries and their use of the application.
Based on an extensive analysis of the diary entries, we categorize and describe use cases for the application in terms of user goal, photo content, location, and photo quality. Common user goals included identifying objects and visual features about them, understanding one’s surroundings, and reading text. We also uncovered some surprising goals, such as avoiding disgusting or dangerous objects, checking for others’ presence, and resolving visual disputes with BLV peers. Participants rated their level of trust and satisfaction with each description, yielding a mean trust score of 2.43 out of 4 (SD=1.16) and a mean satisfaction score of 2.76 out of 5 (SD=1.49). Interestingly, satisfaction and trust scores were not consistently aligned with description accuracy because users were sometimes able to infer information from poor or partially correct descriptions.
We discuss opportunities for future AI-powered scene description applications, arguing that designers should leverage BLV users’ contextual knowledge and experience deciphering imperfect descriptions. We also discuss the importance of exploring AI-powered scene description separately from human-powered visual interpretation and offer future directions for scene description applications.
In summary, we contribute an extensive discussion of use cases for AI-powered scene description applications, along with user assessments of AI-generated descriptions. We hope our work will support researchers and developers as they improve AI-powered accessibility tools, from developing the AI models to designing the applications themselves. In light of the rapid advancement of AI, our work comes at a pivotal moment where we can make AI-powered accessibility tools useful and accessible as they emerge.
2. Related Work
Our work contributes to research on visual descriptions for BLV people and their daily visual descriptive needs. We draw from two related threads: systems that provide visual information about one’s surroundings, and systems that describe images encountered on the Web. We refer to the former category as visual interpretation (human-powered) and scene description (AI-powered) systems, and present work on both approaches (Sections 2.1 and 2.2). We then provide an overview of the latter category, which we refer to as image description systems.
2.1. Human-Powered Visual Interpretation Systems
Researchers have developed human-powered visual interpretation systems that connect BLV users to sighted human assistants. BLV users can use these systems to submit queries with photos to a sighted assistant who then returns an answer within several seconds (Bigham et al., 2010; Naik et al., 2023; Lasecki et al., 2014).
One of the first systems of this kind was VizWiz, a crowd-powered visual question answering (VQA) mobile application deployed by researchers (Bigham et al., 2010). VizWiz received thousands of interpretation requests, yielding a dataset of photos and questions that has been studied extensively (Brady et al., 2013; Gurari et al., 2018, 2019; Tseng et al., 2022). Brady et al. examined submissions from over 5,000 VizWiz users to reveal common visual challenges that blind people encounter in daily life. They created a taxonomy of visual questions and identified situations where the application failed to meet their needs. They found several categories of common questions which were grouped into four high-level categories: identification, description, reading, and unanswerable questions (Brady et al., 2013). Following Brady et al., Gurari et al. found that several mainstream VQA models, despite performing well on the MSCOCO dataset (Lin et al., 2014) and other popular benchmark VQA datasets (Thomee et al., 2016; Deng et al., 2009; Silberman et al., 2012), performed poorly on the VizWiz dataset. Gurari et al. presented this study to encourage the development of CV algorithms that work with BLV users’ photos (Gurari et al., 2018), due to the challenges they present to VQA.
More recent visual interpretation services enable real-time conversations between sighted assistants and users. Unlike VizWiz, sighted assistants can access a video feed of the user’s environment, rather than still images. Formative research gathered user and assistants insights (Avila et al., 2016; Lee et al., 2018; Lannan, 2019). For example, Lee et al. investigated the uses of Aira, a well-known visual interpretation tool, throught two interview studies with sighted assistants and BLV users. In these studies Lee et al. found four use cases for human-powered live visual interpretation: scene description, navigation, task performance (performing tasks), and social engagement (Lee et al., 2018, 2020). Other researchers also investigated ways to improve information quality by studying both trained and untrained visual interpreters’ ability to collaborate to assist one user (Xie et al., 2022, 2023). Their subsequent taxonomy of visual interpretation tasks denote how labor can be divided among multiple agents (Xie et al., 2023).
The above work demonstrates that human-powered systems are useful and enhance BLV people’s quality of life. Investigations into tools like Aira have explored how AI-powered assistance can enhance human support (Lee et al., 2022), but there lacks an examination into the uses of AI-only visual interpretation services. Our work addresses this gap by identifying BLV users’ AI-powered scene description use cases in daily life.
2.2. AI-Powered Scene Description Systems
Researchers have examined AI-powered systems that use CV models to generate visual descriptions of images BLV users upload.
Some researchers have performed studies to record how well AI-powered scene descriptions work in daily life. For example, Smith and Smith performed a one-day autoethnographic diary study to explore the authors’ experiences and perspectives on the usefulness of “readily available [not specified],” AI-powered technologies. In one author’s diary, a visually-impaired researcher, she stated she felt the description application she relied on was most useful when it provided relevant information, and most frustrating when it provided nothing relevant, forcing her to rely on sighted assistance for clarification (Smith and Smith, 2021).
Past work has also evaluated commercialized AI-powered tools, such as SeeingAI (Granquist et al., 2021; Kupferstein et al., 2020). Closest to our work, Kupferstein et al (Kupferstein et al., 2020) presented preliminary results from a pilot diary study exploring four participants’ use of SeeingAI. Participants used SeeingAI when seeking information they wanted to keep private, and participants were frequently unsatisfied with resulting descriptions (only 43% of reactions were positive). This extended abstract revealed that AI can supply information BLV people do not feel comfortable requesting from humans, but it is not clear as to why and when.
Overall, prior research has found that AI-powered applications can address some visual needs of BLV users. However, evaluations have been limited to anecdotal accounts and brief studies. An in-depth approach is necessary to fully understand user experiences and the uses of AI-powered scene description. Our work addresses this with a rigorous diary study that captures natural use of an AI-powered scene description application.
2.3. Image Description
Many researchers have examined descriptions of visual Web content, which are often called “alt-text” since they are communicated through the respective HTML attribute. Image description systems are similar to visual interpretation systems; however, this research mostly concerns images taken by sighted people that are consumed by BLV people browsing the Web.
Past work has explored factors contributing to an image description’s effectiveness in providing relevant information to BLV users (Gleason et al., 2019a; Muehlbradt and Kane, 2022; Stangl et al., 2020, 2021). Stangl et al. conducted interviews with 28 BLV screen reader users, discussing their thoughts on alt-text they received across different types of websites (e.g., news sites, shopping sites, blogs). They found users’ preferences for type and specificity of details in descriptions varied depending on the image’s digital context (Stangl et al., 2020). Researchers have also conducted studies to evaluate the effectiveness of image descriptions (Gleason et al., 2019b; Kreiss et al., 2022; Vedantam et al., 2015; Anderson et al., 2016). For example, Kreiss et al. used participants’ ratings to evaluate the quality of alt-text descriptions in images found in online articles. They recruited 74 participants through Amazon Mechanical Turk, both sighted and BLV, to rate alt-text for images included in an article. They found these ratings could be used as indicators of description quality, and to assess how contextually pertinent information impacted descriptions’ perceived quality (Kreiss et al., 2022).
In summary, no standard has emerged for evaluating descriptions of visual content. Despite this, it is important to establish a methodology for rigorous evaluation that assesses how well a visual description addresses BLV users’ needs. Thus, we took lessons learned from the space of image description and evaluation and applied them to our work. We implemented quality metrics with user-based scores for satisfaction and trust in descriptions and researcher scores on a description’s accuracy. The open-ended field in each diary entry and our follow-up interviews allowed us to collect pertinent yet context-specific information to understand their use cases for AI-powered visual interpretation and the reasoning behind their scores.
3. Methods
Table 1. Participant demographics. The description of vision is self-reported by participants.
PID Age Gender Vision From Birth Vision Description Use SeeingAI Assistive Technology
1
24
Female
Low Vision
N/A
Read with right eye. See colors
Yes
VoiceOver, magnification, and contrast
2
67
Male
Blind
No
Light perception. In extreme conditions identify color.
Yes
VoiceOver
3
57
Male
Blind
Yes
Read with magnification with left eye only.
Yes
CCTV to magnify print, talking thermometer and scale, SeeingAI, and VoiceOver
4
34
Female
Blind
Yes
Quite near-sighted. Sensitivity to light. Able to read with magnification 300%.
Yes
CCTV to magnify print, magnification, VoiceOver, and JAWS
5
55
Male
Blind
No
Only light perception on the periphery
Yes
JAWS,NVDA, VoiceOver, and Voice assistants
6
52
Female
Blind
Yes
No Vision
Yes
VoiceOver, and Voice assistants
7
39
Female
Low Vision
No
Light, color and shape perception. Varies a lot
Yes
ZoomText, and CCTV to magnify
8
44
Male
Blind
Yes
No Vision
Yes
VoiceOver
9
37
Female
Blind
No
No Vision
Yes
JAWS, VoiceOver, and Braille Display
10
39
Male
Blind
Yes
Minimal light perception.
Yes
VoiceOver
11
36
Female
Blind
Yes
No Vision
Yes
VoiceOver, JAWS, and Braille Display
12
32
Female
Low Vision
No
Can see large objects, and colors contrast. Can not read
Yes
JAWS, and VoiceOver
13
39
Female
Blind
No
No Vision
Yes
JAWS, and VoiceOver
14
67
Female
Low Vision
No
Read with magnification.
Yes
JAWS, Braille Display, and magnification
15
47
Female
Blind
No
Light perception, color perception, and shapes perception.
Yes
VoiceOver, and JAWS
16
59
Female
Blind
No
No Vision
Yes
VoiceOver, Voice Assistants, and Braille Display
To examine use cases for scene description applications, we conducted a diary study with 16 BLV users. We developed an application that enabled participants to take photos of their surroundings and receive descriptions of scenes captured in the photos. Participants used this application throughout a two-week period, providing information about their experience every time they received a description.
3.1. Participants
We recruited US-based adult participants through mailing lists from the National Federation of the Blind. We had three inclusion criteria: participants had to self-report non-correctable visual impairment, be over the age of 18, and be willing to use their iPhone, running iOS 9 or higher (application requirement), during the study.
We had 16 participants in total: five participants were men and 11 were women, with ages ranging from 24 to 67 years old (mean=45.5, SD=12.8). Twelve participants identified as blind and 4 identified as low vision. In Table 1, we list these and other self-reported demographics in detail. Participants were compensated with a \$60 Amazon gift card upon completion of the study.
3.2. Procedure
The study consisted of a training session, a two-week application use period, and a follow-up interview. The training session was conducted over a video-conferencing platform and lasted 60 to 90 minutes. During the session, we asked participants about their demographics and experience with AI-powered scene description applications. We then trained them on the use of the study application. This included installing the application on the participant’s mobile device, explaining how to use the application, and submitting a test diary entry (which we did not include in our analysis).
During the diary period, we asked participants to use the application whenever they sought visual information about their surroundings, with a minimum requirement to submit one photo per day. However, we understood that the application, being designed only for scene description, would likely not fulfill every visual need participants had in two weeks (e.g., participants could not read text with the application). Thus, participants were permitted to use other tools to meet their daily needs during the study.
A diary entry included the photo taken, the returned description, and the user’s responses to a brief questionnaire, which was incorporated into the flow of the application. Based on Carroll’s definition of a use case (Carroll, 2003), our questionnaire collected information about the user’s goal and context with the following questions:
  • Where were you when you took the photo? We instructed participants to briefly describe where they used the application (e.g., at my office, at my friend’s apartment, inside the airport, etc.).
  • What information did you want to get from the description? We instructed participants to provide the reason for which they requested visual assistance (e.g., I wanted to know what was outside my window, I was looking for a pair of jeans on my shelf, etc.).
  • How satisfied were you with the description? Answer choices included: Very dissatisfied, Somewhat dissatisfied, Not satisfied nor dissatisfied, Somewhat satisfied, and Very satisfied
  • How much do you trust the description? Answer choices included: Not at all, A little, Somewhat, and To a great extent.
  • Add any additional comments.
We chose different scales for trust and satisfaction, with different numbers and types of options (e.g., “negative” or “positive” for trust, and “negative” to “neutral” to “positive” for satisfaction). For the trust scale, we aimed to capture whether participants were willing to take action based on a description. According to Dumouchel, trust is an action where a person gives something power over themselves, rather than a feeling motivating them; in other words, trust is a choice of action or inaction based on another agent (Dumouchel, 2005). Thus, we chose a 4-point Likert scale for trust because we wanted participants to express their willingness to take action, from “Not at all” (would not take action) to “To a great extent” (would very likely take action). For the satisfaction scale, we aimed to capture whether participants’ expectations were met. Thus, we chose a 5-point scale from “very dissatisfied” to “very satisfied”, with a neutral value in the middle.
After participants completed the diary portion of the study, we conducted semi-structured interviews that lasted 50 to 90 minutes. The interviews were divided into two parts: (1) clarifying use cases from the diary study, and (2) reflecting on the usefulness of the descriptions. During the first part, we inquired about specific diary entries, asking questions like, “What did you do with the information the app provided?” or “Did you want to get a specific piece of information about the background?”. In the second part, we delved into participants’ broader experiences with the application, posing questions such as, “In the context of AI-powered applications, what does trust in an application mean to you?” or “What does satisfaction in an application mean to you?” We asked follow-up questions as needed. Finally, we asked the following to gain insight into participants’ overall evaluation of the application:
  • Overall, how trustworthy do you consider the descriptions provided by the application? Answer choices included: Extremely untrustworthy, Untrustworthy, Somewhat untrustworthy, Neither trustworthy nor untrustworthy, Somewhat trustworthy, Trustworthy, and Extremely trustworthy.
  • Overall, how satisfied are you with the descriptions provided by the application? Answer choices included: Extremely dissatisfied, Dissatisfied, Somewhat dissatisfied, Neither dissatisfied nor satisfied, Somewhat satisfied, Satisfied, and Extremely satisfied.
Figure 2. The scene description application we used to collect data. Screenshots show the flow of using the application and submitting a diary entry. It includes five screens: the photo submission, the photo description, and three diary entry question screens (see 3.2 for questions in questionnaire). The interface was designed to group similar questions, while minimizing the number of elements on each screen.
The scene description application we used to collect data. The screenshots show the flow of using the application and submitting a diary entry. It includes five screens: the photo submission, the photo description, and three diary entry question screens. The photo submission screen has a label that says Eyevisor, a button that says pick photo from album, and a dog in the background that is being captured with the camera. The visual interpretation screen contains a preview of the image captured, the visual interpretation result that says “a dog lying on the floor”, and two buttons “next”, and “back”. The first question screen contained the questions: (1) Where were you when you took the photo and (2) What information did you want to get from the description. The second question screen contained the questions: (3) How satisfied are you with the description, and (4) How much do you trust the description. Finally, in the third questions screen participants could enter additional comments. The interface was designed to group similar questions, while minimizing the number of elements on each screen.
Figure 2. The scene description application we used to collect data. Screenshots show the flow of using the application and submitting a diary entry. It includes five screens: the photo submission, the photo description, and three diary entry question screens (see 3.2 for questions in questionnaire). The interface was designed to group similar questions, while minimizing the number of elements on each screen.
3.3. Scene Description Application
We wanted to create an accessible diary study experience where participants could respond to questions immediately after receiving their photo description (Rieman, 1993). Thus, we developed a custom application to collect our study data. To ensure ecological validity, our application required comparable functionality to commercial applications like Microsoft’s SeeingAI and Google’s Lookout (commonly used by BLV users). To that end, we used Microsoft’s Azure AI Vision image description API since it “exposes the latest deep learning algorithms” (Microsoft, 2023b, a) and is also used by SeeingAI (Team, [n. d.]). This API offers image analysis capable of detecting, classifying objects, and generating descriptions, based on a Microsoft dataset of more than “10,000 [image analysis] concepts” (Microsoft, 2023c, e). Employing this API, our application emulated SeeingAI’s “scene channel,” meaning it could describe environments, identify objects, identify text (but not read it), describe humans, animals, and plants and their actions, and also describe relationships between them.
The flow of the application was simple: when participants launched the application, they could take a photo or upload one from their device’s gallery. Upon submission of a photo, the application presented a results screen that included the description. If satisfied with the results, participants could press the “next” button to load the diary questionnaire (see figure 2). There were three questionnaire screens, and they included five questions in total. In the last questionnaire screen, participants finalized their submission by pressing the “submit” button.
Table 2. Number of diary entries submitted by participants.
Participant ID Number of Diary Entries
1 21
2 32
3 26
4 15
5 21
6 23
7 20
8 20
9 20
10 39
11 23
12 23
13 13
14 21
15 18
16 19
3.4. Qualitative Analysis
We collected a total of 354 diary entries from our 16 participants. On average, each participant submitted 22 entries. The participant with the highest number of submitted entries was 39, and the participant with the lowest number was 13 (see table 2). Out of 354 entries, 22 (6.2%) did not return interpretations due to connectivity or technical issues, so they were discarded from our analysis. Out of the remaining 332, 16 were removed because they were taken by each participant once at the beginning of the study to test the application. The final number of entries analyzed was 316.
The goal of our qualitative analysis was to capture the full context of each diary entry to identify participants’ use cases for scene descriptions. We supplemented our analysis by incorporating user’s perspectives from the follow-up interviews.
Our analysis involved two parts: (1) an inductive open coding process of the diary entries and the follow-up interviews (Saldaña, 2021), and (2) affinity diagramming sessions to group the resultant codes (Hartson and Pyla, 2012). For the diary entries, we coded participants’ questionnaire responses and the content of their submitted photos. For the follow-up interviews, the open coding process organically revealed connections with diary codes, and also revealed underlying themes behind the use cases for scene description applications. Finally, in the affinity diagramming sessions, we grouped codes to create the themes we share in Findings (see Section 4).
For the diary entries, we derived codes based on the participant’s goal in the entry and context in which they used the application. We defined User Goal as the reason for taking the photo (e.g., identifying a specific object, determining an object’s color), drawn from participant’s self-reported goals (diary entry response to “What did you want to get from the description?”), photo content, and statements in supporting interviews. Most diary entries had one goal, but some had multiple. To code context, we examined all information relating to the “what was captured” and “where it happened” of each diary entry. The “what” included the content of the photo and photo properties based on visual qualities that may impact the API’s performance (e.g., lighting conditions, photo clarity). The “where” included the self-reported location where the photo was taken and the image’s setting (e.g., indoors, outdoors, private, public). Two researchers coded 30 diary entries individually, then came together to discuss discrepancies and generated a codebook. Afterwards, remaining entries were split between researchers and codes were added or clarified as needed.
Table 3. Shows the five coded user goal levels, definitions, and examples applied under each level.
User Goal Definition User Goal Examples Image Example
Describe Scene
Gathering visual information about related subjects that form a scene
“Is there a mountain in the background?”
“Was my room messy or not?”
Learn Application & Learning how to use the application, testing how well the application performed, or becoming familiar with the application output “I was testing to see if it could describe the pattern on this shirt.” “I wanted to use the app to help me frame a photo of my dog.” Identify Subject Identifying a specific subject when the user has no clear working knowledge of the subject “What was the image my friend just tweeted?” “What item am I holding?” Identify Feature Identifying features of a subject such as color, size, state (on/off, full/empty), and so on, when a user already has a working knowledge of the subject’s identity “Is the lamp in my room on or off?”
“What is the logo on this baseball glove?” No Specific Goal Goal was either unclear to the coders, or the user mentioned they had no goal in mind “I just wanted a basic description”
“People”
Table 4. Part 1 of coded locations. Shows four coded location levels, definitions, and examples applied under each level.
Location Location Examples Image Example
Used in educational or spiritual sites
Schools, churches, and
other religious institutions
[…]
This photo represents an example of an entry coded for location “Used in healthcare sites”. Participant was located in an optometrist office. A photo of a group of medical equipments for general health checks and an examination table.
Used in recreational or leisure areas
Restaurants, parks, and historical sites
[…]
This photo represents an example of an entry coded for location “Used in recreational or leisure areas”. Participant was located in a historical site. A photo of a woman taking a selfie in an open field with an aztec pyramid in the background.
Used in healthcare places
Hospitals, dentist offices, and ophthalmologist offices
[…]
This photo represents an example of an entry coded for location “Used in educational or spiritual sites”. Participant was located in a chapel. A photo of an empty chapel with wooden pews.
Used in travel services sites
Bus stops, train stations, and airports
[…]
This photo represents an example of an entry coded for location “Used in travel service sites”. Participant was located in an airport. A photo of a gateway waiting area at an airport full of people.
Table 5. Part 2 of coded locations. Shows the other four coded location levels, definitions, and examples applied under each level.
Location Location Examples Image Example
Used in retail businesses
Supermarkets, hair salons, and other businesses
[…]
This photo represents an example of an entry coded for location “Used in retail businesses”. Participant was located in a supermarket, at the alcohol section. Shelves full of wine bottles at a supermarket
Used at participant’s work
Any spaces participants defined as their “work”, such as corporate or home offices
[…]
This photo represents an example of an entry coded for location “Used at participant’s work”. Participant was in their home studioin recording their podcast. A photo of a black and silve microphone on a stand with several electronic devices on the table
Used in living spaces
Apartments, senior homes, and houses
[…]
This photo represents an example of an entry coded for location “Used in living spaces or domiciles”. Participant was located at an assisted living facility. A photo of a stripe patterned drinking cup on a table. In the background there are people on wheel chairs.
Used in unidentified locations
Applied when neither the picture nor the participant’s diary responses could tell us where they were located
[…]
This photo represents an example of an entry coded for location “Used in unidentified areas”. Participant was located in a indoor space setting but in their dairy entry they did not mention where they were. A photo of a payphone and a trash can.
To code follow-up interviews, audio recordings were transcribed by an automatic transcription service and corrected by one researcher. Then, two researchers coded four interviews individually. Finally, the researchers discussed disparities to add interview-derived codes to the codebook, and remaining interviews were split between researchers for coding.
These processes resulted in over 200 codes in total, which we grouped using multiple rounds of affinity diagramming. One researcher established three initial categories for code groupings: (1) Properties of the photos, (2) Content of the photos, and (3) Uses of the application. These were refined into subcategories (e.g., subcategories for objects captured as photo content such as living subjects, recreational objects and decor, etc.). Afterwards, we conducted thematic analysis using two rounds of affinity diagramming, one that examined grouped codes solely from the diary entries, and a second that incorporated interview codes. This generated our final themes (see Section 4.1 and 4.2).
After coding and creating themes, two researchers also scored the accuracy of each entry’s scene description. We wanted to determine the appropriateness of the application descriptions given what information was available in submitted photos. This accuracy score allowed us to: (1) understand whether a scene description contained “usable” information, (2) understand how frequently the scene description application provided “reasonable” descriptions, and (3) find relationships between participants’ reactions and the accuracy of the descriptions. We defined accuracy as, “the extent to which the scene description matched a sighted human perception of the image,” adapted from user-based evaluations of image descriptions (Kreiss et al., 2022). Two researchers coded two participants’ diaries individually, then came together and split remaining entries, scoring each image-interpretation pair on a scale of 1 to 3, where 1 meant the interpretation did not match a human perception at all, 2 meant a partial match, and 3 meant it mostly matched. We share results about accuracy scores and discuss how participants’ definitions of satisfaction varied based on description accuracy (see Section 4.3).